(i) Let V be a Banach space. Prove that if V* separable then V is not separable.


(ii) Give an example of separable Banach space V which has a non-separable dual space V* . 


1
Expert's answer
2021-02-24T06:10:11-0500

i) Proving that aBanach space is separable if its dual is separable:

Let "\{f_n\}_{n=1}^{\infin}" be a dense subset of the unit ball "B" in "X^*". For each "n\isin N", pick "x_n\isin X" such that "f_n(x_n)>\frac{1}{2}" . Let "Y=\overline{span\{x_n\}}" and observe that "Y" is separable, since finite rational combinations of "\{x_n\}" are dense in "Y". It is now sufficient to show that "X=Y". We proceed by contradiction. Suppose that "X\neq Y". Then there is an "f\isin X^*", with "||f||=1" such that "f(x)=0" for all "x\isin Y" . Now choose "n" such that "||f_n-f||<\frac{1}{4}". Then

"|f(x_n)|=|f_n(x_n)-f_n(x_n)+f(x_n)|\geq |f_n(x_n)|-|f_n(x_n)-f(x_n)|\geq"

"\geq |f_n(x_n)|-||f_n-f||\cdot ||x_n)||>\frac{1}{2}-\frac{1}{4}=\frac{1}{4}"


ii) Space "L^p"